OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –3 5 0 4 Line 2 x y 0 4 6 2 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

@jdoe0001 @just_one_last_goodbye @k_lynn @bibby @Bookworm14 @HelpBlahBlahBlah @BassCatcher15 @xapproachesinfinity @xX_GamerGirl_Xx @Squirrels @HelpBlahBlahBlah @Godlovesme @EclipsedStar @EmmaElizabeth @abb0t

OpenStudy (jdoe0001):

well, you could start off by getting their slopes if their slopes are equal, they're parallel lines, thus they do not intersect, thus no solution if not, then they do touch, thus 1 solution

OpenStudy (anonymous):

So A? @jdoe0001

OpenStudy (jdoe0001):

hmmm well.. what slopes did you get?

OpenStudy (anonymous):

for line 1 -3,4

OpenStudy (jdoe0001):

3,4? may want to revise your slopes

OpenStudy (jdoe0001):

\(\large { \begin{array}{ccllll} x&y \\\hline\\ {\color{red}{ -3}}&{\color{blue}{ 5}}\\ {\color{red}{ 0}}&{\color{blue}{ 4}} \end{array}\qquad \begin{array}{ccllll} x&y \\\hline\\ {\color{red}{ 0}}&{\color{blue}{ 2}}\\ {\color{red}{ 4}}&{\color{blue}{ 6}} \end{array}\qquad slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} }\)

OpenStudy (anonymous):

-3y-0y/0x-4x

OpenStudy (anonymous):

or is it -3y-4y/2x-6x

OpenStudy (jdoe0001):

hmm shoot I got the numbers off

OpenStudy (anonymous):

its okay

OpenStudy (jdoe0001):

anyhow... you'd need to get the slopes first for each \(\large { \begin{array}{ccllll} x&y \\\hline\\ {\color{red}{ -3}}&{\color{blue}{ 5}}\\ {\color{red}{ 0}}&{\color{blue}{ 4}} \end{array}\qquad \begin{array}{ccllll} x&y \\\hline\\ {\color{red}{ 0}}&{\color{blue}{ 4}}\\ {\color{red}{ 6}}&{\color{blue}{ 2}} \end{array}\qquad slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} }\)

OpenStudy (jdoe0001):

if the slopes are equal, no solution, or infinite solutions if they differ, then 1 solution

OpenStudy (anonymous):

i understand that i just dont understand slope

OpenStudy (jdoe0001):

well... if you have not covered slopes yet... how are you meant to find whether they have solution or not then? that is, what is the chapter section covering to do so?